If 6.55 g of CuNO3 is dissolved in water to make a 0.690 M solution, what is the volume of the solution?
Answer #1

molar mass of CuNO3 = 63.5 +14+3*16 = 125.5g

1L of 0.69M contins 0.69 mole so a mass of 0.69*125.5 = 86.59 g

and 6.55 g corresponds to 6.55/86.59 = 0.0756 L = 75.6mL